OpenStudy (anonymous):

tanx = 1/3 and tany = 1/7 then 2x + y is?

OpenStudy (anonymous):

2x = tan^-1 2/3 y = tan^-1 1/7

hartnn (hartnn):

can u find tan 2x first ?

hartnn (hartnn):

no, not like that tan 2x = (2tanx)/(1+tan^2 x)

OpenStudy (anonymous):

(2tanx)/(1-tan^2 x).........

hartnn (hartnn):

oh,yes..... when u are done finding tan 2x use tan (A+B)= ..... u know

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

tan 2a = 3/4

OpenStudy (anonymous):

@Yahoo! u were right

OpenStudy (anonymous):

2x = tan^-1 3/4 y = tan^-1 1/7 2x + y = (3/4 + 1/7)/1-3/21

OpenStudy (anonymous):

25/17

hartnn (hartnn):

tan 2a is correct... u got 25/17 as final answer ?

OpenStudy (anonymous):

yes......

OpenStudy (anonymous):

~~~~ tan pie/3

hartnn (hartnn):

isn't the answer 1 ?

OpenStudy (anonymous):

is 25/17 correct

OpenStudy (anonymous):

Oh..no...it is 25/25

OpenStudy (anonymous):

Yes....it is pie/4

OpenStudy (anonymous):

thxxx

hartnn (hartnn):

welcome :)